# Anomalous bulk-edge correspondence of nonlinear Rice-Mele model

### v1

@@ III.2 Eigenvalue’s nonlinearity and the corresponding bulk-edge correspondence
In the preceding subsection (III.1), we presented an overview of four distinct categories of eigenvalue problems, which are concisely summarized in Table 1. In the subsequent subsection (III.2), we adopt the methodology detailed in Ref. [64] to delve into the anomalous nonlinearity of the eigenvalues within the nonlinear RM Hamiltonian (1) and its corresponding BEC. Our strategy for examining this anomalous eigenvalue nonlinearity in the nonlinear RM Hamiltonian (1) and BEC, given the underlying nonlinearity of the eigenvalues, involves the utilization of auxiliary eigenvalues, as outlined below:
@@
$H_{\text{non}}\left(\omega,t\right)\varPsi=\omega S\left(\omega,t\right)\varPsi,$ (4)
In Equation (4), the time-dependent $H\left(\omega,t\right)$ represents the Hermitian matrix corresponding to the nonlinear RM Hamiltonian (1). The symbol $\varPsi$ denotes the nonlinear eigenfunction. Moreover, the $S\left(\omega,t\right)$ is the overlap matrix, and $\omega$ serves as the parameter that characterizes the anomalous nonlinearity compared with the intrinsic nonlinearity of $g$ in the nonlinear RM Hamiltonian (1) .
@@
From Equation (5), the solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (4). In order to gain a more profound understanding of the BEC model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$.
@@
with $S_{0}$ being a diagonal matrix given by:
@@
Next, we proceed to investigate the impact of the interaction parameter $g$ on BEC by studying the evolution of the position expectation value $\langle X\rangle=\sum_{j}j|\Psi_{j}|^{2}$ for the soliton state over a single adiabatic cycle. To achieve this, we numerically solve Eqs. (2) and (3), employing two distinct methodologies. In the first approach, we determine the soliton at time $t$ through an iterative process, aiming to reach the steady state. The starting point for each iteration is the instantaneous soliton obtained from the previous time step. In the second method, we obtain the soliton at time $t$ by directly applying a given initial soliton and solving the equation for the time evolution of the wavefunction using the fourth-order Runge-Kutta method. It is important to highlight that when the eigenvalues of the RM model transition to nonlinear eigenvalues, the pumping behavior of the system exhibits significant changes. Specifically, at $g=0$, the pump transport oscillates near the initial point, as shown in Fig. 2(a2). As the interaction strength increases to $g=1$ and $g=5$, the system demonstrates pump transport phenomena. It is essential to note that when $g=1$, the pump transport value for one period deviates from 2, as depicted in Fig. 2(b2). Furthermore, when the interaction strength continues to rise to $g=5$, the soliton wave dynamical evolve within a period decreases more pronounced, as depicted in Fig. 2(c2). Consistent with the original RM model, when $g=10$, the pumping ceases to operate in the RM model with the nonlinear eigenvalue problem, as illustrated in Fig. 2(d2).
@@
Finally, note that this section provides a thorough examination of the eigenvalue problem in Hamiltonian systems, contrasting the linear and nonlinear cases. Linear Hamiltonians serve as the cornerstone for determining energy states in both classical and quantum mechanics. In contrast, nonlinear Hamiltonians, which depend on the magnitude of the wave function, introduce more intricate challenges. We also discuss how the incorporation of eigenvalue terms that depend on state variables enriches linear Hamiltonian systems, giving rise to phenomena such as level crossing. Furthermore, we extend the concept of nonlinear eigenvalues to nonlinear Hamiltonian systems, emphasizing their vital role in modeling complex quantum states described by non-harmonic potentials. By introducing an auxiliary eigenvalue $\lambda$, we derive the time-energy spectrum of $\lambda$ and investigate the temporal evolution of the expectation value of the position for the ground state. Our findings reveal that the inclusion of nonlinear eigenvalues primarily influences the pump transmission and stability of the ground state solitons.
@@ IV Nonlinear bulk-edge correspondence
Figure 3: Effects of parameter $\Delta$ and $g$ on the auxiliary $\lambda$-spectrum defined in Eq. (6) for the nonlinear RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $T=2000\pi$ and $\omega=10^{-3}$. Other parameters are given as (a1): $g=0$, $\Delta=0.5$; (a2): $g=0$, $\Delta=0$; (b1): $g=10$, $\Delta=0.5$; (b2): $g=10$, $\Delta=0$.
In the preceding Sec. III, we examined four distinct eigenvalue problems and computed the anomalous nonlinear eigenvalue problem for the nonlinear RM Hamiltonian (1), obtaining its $\lambda$-spectrum and the pumping diagram for the ground state. Following this, Sec. IV presents a comprehensive analysis of how the parameter $\Delta$ influences the anomalous nonlinear eigenvalue problem in the nonlinear RM Hamiltonian (1).
@@
Figure 4: (a)-(c): Characterize the relationship between time $t$ and frequency $\omega$ at $\lambda=0$ for the RM model with $g=2$. Parameter $\Delta$ are set to $\Delta=1$, $\Delta=0.5$, and $\Delta=0$, respectively. (d): The Chern numbers calculated for each $\Delta$ and $\omega$, red indicates a Chern number $\mathcal{C}=1$; blue indicates $\mathcal{C}=0$; green indicates $\mathcal{C}=-1$.
@@
We proceed to plot the $t-\omega$ diagrams for different values of $\Delta$ ($\Delta=1$, $\Delta=0.5$, and $\Delta=0$), as shown in Figs. 4(a)-(c) under the conditions of $\lambda=0$ and $g=2$. When $\Delta=1$, the edge states exhibit a pronounced turning point, as illustrated in Fig.4(a), which is in accordance with the edge state crossing observed in Fig.2(a1). In contrast, for $\Delta=0.5$, the edge states cease to display any turning behavior, as depicted in Fig.4(b). This observation aligns with the convergence of edge states at the conclusion of the time period shown in Fig.3(a1) and (a2), hinting at the occurrence of a nonlinear topological phase transition. Lastly, when $\Delta=0$, the edge states remain uncrossed and retain a parallel configuration, as demonstrated in Fig.4(c). This result is consistent with the parallel edge states observed in Fig.3(b1) and (b2).
To elucidate the observed variations in the edge and soliton states presented in the aforementioned figures, we employ the topological invariant of the system, namely the Chern number associated with the RM model. The Chern number for the first non-degenerate band of a two-dimensional system is mathematically defined as the integral of the Berry curvature over the entire Brillouin zone, which can be expressed as follows [1, 64]
[…]
@@
where $\mathcal{A}_{1}(\vec{R})=i\langle\Psi_{1}(\vec{k},t)|\nabla_{\vec{k}}|\Psi_{1}(\vec{k},t)\rangle$ is the Berry connection for the ground state, $|\Psi_{1}(\vec{k},t)\rangle$ is the auxiliary eigenstate of the ground state, and $\vec{R}$ represents the integration boundaries, specifically $\vec{R}=\left[0;2\pi\right]\times\left[0;T\right]$. In the RM model at $g=0$, with the parameter $\omega$ fixed at $10^{-3}$, it is observed that the Chern number adopts non-zero values when $\left|\Delta\right|>0.5$. Our numerical computations reveal that the ground state band presented in Fig. 2(a1) exhibits a Chern number of $\mathcal{C}_{1}=1$. In contrast, the ground state as shown in Fig. 3(b1) shows a Chern number of $\mathcal{C}_{1}=0$. In particular, the ground state as showed in Fig. 3(a1) shows the critical points of Chern number. These results strongly indicate a profound correlation between the value of the Chern number and the count of points where $\lambda=0$.
@@
To delve deeper into the connection between the Chern number and the band structure of the anomalous eigenvalue’s nonlinearity of nonlinear RM model, we construct a phase diagram in the $\Delta-\omega$ parameter space, as shown in Fig. 4(d). This phase diagram encapsulates two distinct phases of the nonlinear RM model: a topologically nontrivial phase characterized by a nonzero Chern number $\mathcal{C}_{1}$ (represented by red and yellow regions), and a topologically trivial phase $\mathcal{C}_{1}=1$, while the yellow region corresponds to $\mathcal{C}_{1}=-1$. By fixing $\omega=10^{-3}$ and drawing the corresponding line on the phase diagram, we can identify the intersection points with the phase boundaries as the locations where changes occur in the edge states. These observations underscore a profound connection between the band structure of the nonlinear RM model and the Chern number.
@@
Therefore, it is concluded that the eigenvalue’s nonlinearity of nonlinear RM model exhibits a remarkable capability to modulate the behavior of edge states within the time-energy spectrum through the adjustment of the parameter $\Delta$. Specifically, when $g=0$, a decrease in $\Delta$ from 1 to 0.5 induces a rightward shift in the intersections of edge states during the latter portion of the period. As $\Delta$ continues to diminish towards 0, these intersections gradually diminish, ultimately culminating in the closure of the band gap at $t=0$ when $\Delta=0$. Similarly, for $g=10$, analogous behavior is observed. However, a distinctive phenomenon emerges at $\Delta=0.3285$, characterized by the appearance of additional bands. This manifestation is evidenced by the emergence of an extra soliton state in proximity to the ground state and the appearance of edge states during the first half of the period, both of which are intricately connected to the Chern number. When $g=0$ and $|\Delta|>0.5$, the Chern number adopts non-zero values, and the model undergoes a nonlinear topological phase transition precisely at $\Delta=0.5$. The phase diagram in the $\Delta-\omega$ parameter space delineates the boundaries separating topologically nontrivial and trivial phases. The intersections of these phases at specific $\omega$ values unveil alterations in the edge states, further underscoring the intricate interplay between the Chern number and the band structure of the nonlinear RM model.
@@ V Conclusion and outlook
This work delves into the behavior of the RM model across varying degrees of nonlinearity, unveiling the intricate interplay between topological attributes and nonlinear dynamical processes. The research unequivocally demonstrates a transition from a linear Hamiltonian regime, typified by a topological insulator harboring protected edge states, to a regime manifesting solitons and nonlinear pumping phenomena upon the introduction of nonlinearity. By incorporating an auxiliary eigenvalue $\lambda$, the study offers a novel vantage point on the time-energy spectrum and the temporal progression of the ground state, underscoring the pivotal role of the nonlinear eigenvalue in governing pumping transport and stability.
@@
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Nan Li, Shujie Cheng, Chenhui Yan, and Biao Wu for stimulating discussions and useful help. This work was supported by the National Natural Science Foundation of China (Grants No. 12074344, No. 12374246) and the Zhejiang Provincial Natural Science Foundation (Grant No. LZ21A040001).

### v2

@@ III.2 Eigenvalue’s nonlinearity and the corresponding bulk-edge correspondence
In Sec. III.1, we presented an overview of four distinct categories of eigenvalue problems, which are concisely summarized in Table 1. In the subsequent Sec. III.2, we adopt the methodology detailed in Ref. [64] to delve into the anomalous nonlinearity of the eigenvalues within the nonlinear RM Hamiltonian (1) and its corresponding bulk-edge correspondence. Our strategy for examining this anomalous eigenvalue nonlinearity in the nonlinear RM Hamiltonian (1) and bulk-edge correspondence, given the underlying nonlinearity of the eigenvalues, involves the utilization of auxiliary eigenvalues, as outlined below.
@@
$H_{\text{non}}\left(\omega,t\right)\varPsi=\omega S\left(\omega,t\right)\varPsi.$ (4)
In Eq. (4), the time-dependent $H\left(\omega,t\right)$ represents the Hermitian matrix corresponding to the nonlinear RM Hamiltonian (1). The symbol $\varPsi$ denotes the nonlinear eigenfunction. Moreover, the $S\left(\omega,t\right)$ is the overlap matrix, and $\omega$ serves as the parameter that characterizes the anomalous nonlinearity compared with the intrinsic nonlinearity of $g$ in the nonlinear RM Hamiltonian (1). It is worth noting that Eq. (4) together with $S$ are employed to introduce auxiliary eigenvalues and furthermore the topological edge states of auxiliary eigenstates are topologically inherited as physical edge states, which represents interplay between the topology and nonlinearity of eigenvalues.
@@
From Eq. (5), the solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (4). To gain a more profound understanding of the bulk-edge correspondence model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$.
@@
with $S_{0}$ being a diagonal matrix given by
@@
Next, we proceed to investigate the impact of the interaction parameter $g$ on bulk-edge correspondence by studying the evolution of the position expectation value $\langle X\rangle=\sum_{j}j|\Psi_{j}|^{2}$ for the soliton state over a single adiabatic cycle. To achieve this, we numerically solve Eqs. (2) and (3), employing two distinct methodologies. In the first approach, we numerically propagate the initial soliton profile through time evolution governed by the nonlinear Schrödinger equations [Eqs. (2) and (3)], implemented via a fourth-order Runge-Kutta algorithm. This method directly simulates the dynamical evolution of the wave function under physical time-dependent driving. In the second approach, we employ an iterative self-consistent scheme to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation. Under adiabatic conditions ($\omega\rightarrow 0$), both methods yield identical results, thereby mutually validating their accuracy. For nonadiabatic driving ($\omega\gg 0$), the first method faithfully reproduces experimentally observable dynamics [e.g., soliton jumps in Figs. 1(c2) and 2(c2)], while the second method serves as a diagnostic tool to quantify deviations from adiabaticity by comparing instantaneous eigenstates with dynamically evolved states. We remark that in the second approach, the wave function evolves through iterative steps as an iterative term, thereby ensuring the successful derivation of the band structure for the nonlinear wave function. It is important to highlight that when the eigenvalues of the RM model transition to nonlinear eigenvalues, the pumping behavior of the system exhibits significant changes. Specifically, at $g=0$, the pump transport oscillates near the initial point, as shown in Fig. 2(a2). As the interaction strength increases to $g=1$ and $g=5$, the system demonstrates pump transport phenomena. When $g=1$, the pump transport value over one period remains 2, as depicted in Fig. 2(b2). Furthermore, when the interaction strength continues to rise to $g=5$, the soliton wave dynamical evolve within a period decreases more pronouncedly, as depicted in Fig. 2(c2). Consistent with the original RM model, when $g=10$, the pumping ceases to operate in the RM model with the nonlinear eigenvalue problem, as illustrated in Fig. 2(d2).
@@
Finally, note that this section provides a thorough examination of the eigenvalue problem in Hamiltonian systems, contrasting the linear and nonlinear cases. Linear Hamiltonians serve as the cornerstone for determining energy states in both classical and quantum mechanics. In contrast, nonlinear Hamiltonians, which depend on the magnitude of the wave function, introduce more intricate challenges. We also discuss how the incorporation of eigenvalue terms that depend on state variables enriches linear Hamiltonian systems, giving rise to phenomena such as level crossing. Furthermore, we extend the concept of nonlinear eigenvalues to nonlinear Hamiltonian systems, emphasizing their vital role in modeling complex quantum states described by nonharmonic potentials. By introducing an auxiliary eigenvalue $\lambda$, we derive the time-energy spectrum of $\lambda$ and investigate the temporal evolution of the expectation value of the position for the ground state. Our findings reveal that the inclusion of nonlinear eigenvalues primarily influences the pump transmission and stability of the ground-state solitons.
@@ IV Nonlinear bulk-edge correspondence
Figure 3: Effects of parameter $\Delta$ and $g$ on the auxiliary $\lambda$ spectrum defined in Eq. (6) for the nonlinear RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $T=2000\pi$, and $\omega=10^{-3}$. Other parameters are given as (a1): $g=0$, $\Delta=0.5$; (a2): $g=0$, $\Delta=0$; (b1): $g=10$, $\Delta=0.5$; (b2): $g=10$, $\Delta=0$.
In the preceding Sec. III, we examined four distinct eigenvalue problems and computed the anomalous nonlinear eigenvalue problem for the nonlinear RM Hamiltonian (1), obtaining its $\lambda$ spectrum and the pumping diagram for the ground state. Following this, Sec. IV presents a comprehensive analysis of how the parameter $\Delta$ influences the anomalous nonlinear eigenvalue problem in the nonlinear RM Hamiltonian (1).
@@
Figure 4: (a)-(c) Characterization of the relationship between time $t$ and frequency $\omega$ at $\lambda=0$ for the RM model with $g=2$. Parameter $\Delta$ are set to $\Delta=1$, $\Delta=0.5$, and $\Delta=0$, respectively. (d) The Chern numbers calculated for each $\Delta$ and $\omega$, red indicates a Chern number $\mathcal{C}=1$; blue indicates $\mathcal{C}=0$; green indicates $\mathcal{C}=-1$.
To investigate the influence of the model parameter $\Delta$ on the nonlinearity of the anomalous eigenvalue associated with the nonlinear RM Hamiltonian [Eq. (1)], with a specific focus on its effect on the $t-\omega$ relationship, we will generate $t-\omega$ diagrams for a range of $\Delta$ values while maintaining $\lambda=0$. Meanwhile, it is worth highlighting that soliton states emerge independently of the specific value of $\Delta$. Moreover, the count of soliton states escalates with the interaction strength g. To facilitate a clearer understanding, we focus our examination on the scenario where $g=2$, which results in the presence of two soliton states, thereby enabling a more straightforward analysis and observation of our findings.
We proceed to plot the $t-\omega$ diagrams for different values of $\Delta$ ($\Delta=1$, $\Delta=0.5$, and $\Delta=0$), as shown in Figs. 4(a) to 4(c) under the conditions of $\lambda=0$ and $g=2$. When $\Delta=1$, the edge states exhibit a pronounced turning point, as illustrated in Fig. 4(a), which is in accordance with the edge state crossing observed in Fig. 2(a1). It is observed that in the band diagram of Fig. 4(a), the edge states exhibit sharp bending due to the emergence of self-crossing structures within the energy bands of the nonlinear Bloch system. In contrast, for $\Delta=0.5$, the edge states cease to display any turning behavior, as depicted in Fig. 4(b). This observation aligns with the convergence of edge states at the conclusion of the time period shown in Figs. 3(a1) and 3(a2), hinting at the occurrence of a nonlinear topological phase transition. Lastly, when $\Delta=0$, the edge states remain uncrossed and retain a parallel configuration, as demonstrated in Fig. 4(c). This result is consistent with the parallel edge states observed in Figs. 3(b1) and 3(b2).
To elucidate the observed variations in the edge and soliton states presented in the aforementioned figures, we employ the topological invariant of the system, namely the Chern number associated with the RM model. The Chern number for the first nondegenerate band of a two-dimensional system is mathematically defined as the integral of the Berry curvature over the entire Brillouin zone, which can be expressed as follows [1, 64]:
@@
where $\mathcal{A}_{1}(\vec{R})=i\langle\Psi_{1}(k,t)|\nabla_{k}|\Psi_{1}(k,t)\rangle$ is the Berry connection for the ground state. Here, $|\Psi_{1}(k,t)\rangle$ is the wave function in the momentum space of $k$ by making a Fourier transformation of the wave function $|\psi_{n}(t)\rangle$ in Eq. (1). We remark that for a one-dimensional dimerized lattice (with two sites per unit cell), the transformation of the discrete real-space wave function $|\Psi_{n}(t)\rangle$ into momentum $k$ space requires careful treatment of the unit cell structure. Consider a system with $N$ unit cells, where each unit cell contains two lattice sites (indexed as $n=1,2,…,2N$). We define the unit cell index $j=1,2,...,N$ and partition the lattice into even-indexed and odd-indexed site sublattices. The momentum-space wave functions for each sublattice are obtained via a discrete Fourier transformation: $|\Psi_{A}(k,t)\rangle=1/\sqrt{N}\sum_{j=1}^{N}e^{ikj}|\Psi_{2j}(t)\rangle$ and $|\Psi_{B}(k,t)\rangle=1/\sqrt{N}\sum_{j=1}^{N}e^{ikj}|\Psi_{2j-1}(t)\rangle$ with $k=2\pi m/N$ ($m=0,1,...,N-1$). The total momentum-space wave function is then expressed as a two-component spinor as $|\Psi_{1}(k,t)\rangle=(|\Psi_{A}(k,t),|\Psi_{B}(k,t))^{T}$. Here since the model system is one-dimensional, $k$ is a scalar. The vector $\vec{R}$ represents the integration boundaries, specifically $\vec{R}=\left[0;2\pi\right]\times\left[0;T\right]$. In the RM model at $g=0$, with the parameter $\omega$ fixed at $10^{-3}$, it is observed that the Chern number adopts nonzero values when $\left|\Delta\right|>0.5$. Our numerical computations reveal that the ground-state band presented in Fig. 2(a1) exhibits a Chern number of $\mathcal{C}_{1}=1$. In contrast, the ground state as shown in Fig. 3(b1) shows a Chern number of $\mathcal{C}_{1}=0$. In particular, the ground state as showed in Fig. 3(a1) shows the critical points of Chern number. These results strongly indicate a profound correlation between the value of the Chern number and the count of points where $\lambda=0$.
@@
To delve deeper into the connection between the Chern number and the band structure of the anomalous eigenvalue’s nonlinearity of nonlinear RM model, we construct a phase diagram in the $\Delta-\omega$ parameter space, as shown in Fig. 4(d). This phase diagram encapsulates two distinct phases of the nonlinear RM model this diagram illustrates a topologically nontrivial phase, identified by a nonzero Chern number $\mathcal{C}_{1}$ (shown in red and green regions), coexisting with a topologically trivial phase where $\mathcal{C}_{1}=0$ (blue regions). Specifically, the green region corresponds to $\mathcal{C}_{1}=-1$. By fixing $\omega=10^{-3}$ and drawing the corresponding line on the phase diagram, we can identify the intersection points with the phase boundaries as the locations where changes occur in the edge states. These observations underscore a profound connection between the band structure of the nonlinear RM model and the Chern number.
@@
Therefore, it is concluded that the eigenvalue’s nonlinearity of nonlinear RM model exhibits a remarkable capability to modulate the behavior of edge states within the time-energy spectrum through the adjustment of the parameter $\Delta$. Specifically, when $g=0$, a decrease in $\Delta$ from 1 to 0.5 induces a rightward shift in the intersections of edge states during the latter portion of the period. As $\Delta$ continues to diminish towards 0, these intersections gradually diminish, ultimately culminating in the closure of the band gap at $t=0$ when $\Delta=0$. Similarly, for $g=10$, analogous behavior is observed. However, a distinctive phenomenon emerges at $\Delta=0.3285$, characterized by the appearance of additional bands. This manifestation is evidenced by the emergence of an extra soliton state in proximity to the ground state and the appearance of edge states during the first half of the period, both of which are intricately connected to the Chern number. When $g=0$ and $|\Delta|>0.5$, the Chern number adopts nonzero values, and the model undergoes a nonlinear topological phase transition precisely at $\Delta=0.5$. The phase diagram in the $\Delta-\omega$ parameter space delineates the boundaries separating topologically nontrivial and trivial phases. The intersections of these phases at specific $\omega$ values unveil alterations in the edge states, further underscoring the intricate interplay between the Chern number and the band structure of the nonlinear RM model.
@@ V Conclusion and outlook
This work delves into the behavior of the RM model across varying degrees of nonlinearity, unveiling the intricate interplay between topological attributes and nonlinear dynamical processes. The research unequivocally demonstrates a transition from a linear Hamiltonian regime, typified by a topological insulator harboring protected edge states, to a regime manifesting solitons and nonlinear pumping phenomena upon the introduction of nonlinearity. By incorporating an auxiliary eigenvalue $\lambda$, the study offers a vantage point on the time-energy spectrum and the temporal progression of the ground state, underscoring the pivotal role of the nonlinear eigenvalue in governing pumping transport and stability.
@@
Acknowledgements
We thank Ying Hu, Yapeng Zhang, Nan Li, Shujie Cheng, Chenhui Yan, and Biao Wu for stimulating discussions and useful help. This research was supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No. LZ25A040004 and the National Natural Science Foundation of China under Grants No. 12074344.
Data availability
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.